# Hyperdimensional Computing for Node Classification and Link Prediction

### v2

@@ 5.6 Results
Table 4: Class Incremental Node Labeling Results. No hyperparameter grid search is used for GCN and SGC algorithms. Wall-time includes both training and inference of the model. HDGL does not require retraining from $t=2$.
@@
It is crucial to note that while we have used grid search on 12 potential setups for the deep learning model hyperparameter optimization of deep learning models typically explores a much larger space. This expanded search naturally comes with increased runtime, especially when GPUs are utilized, further widening the energy efficiency gap between Graph Neural Networks (GNN) and traditional deep learning models like HDGL.
@@
Figure 4: Link Prediction Performance of HDGL under Various dimensionality configurations.
HDGL in the Class-Incremental Learning Setting Consider a scenario with a dynamic graph where nodes and edges remain static, but the new node labels appear over time i.e initially, only a subset of nodes is labeled, belonging to two classes. Over time, new nodes receive labels; thus, introducing new classes. This is a node labeling instance of the class-incremental learning problem [52, 3, 51]. In this setting, deep learning methods typically require either retraining on the data for previously learned classes along with the data for the new classes to avoid catastrophic forgetting of the previously learned classes [51]. This retraining incurs significant additional runtime and computation.
@@
To simulate incremental node classification on a fully labeled graph, we use BlogCatalog and Coauthor Physics Dataset. We first split the nodes into train/val/test as in Shchur et al. [45]. We mask the train/val/test such that at time $t=1$ only the labels $\{1,2\}$ are available; and at $t=2$ only the labels $\{1,2,3\}$ are only available; and so on such that at the final time step, all of the labels $\{1,\cdots,L\}$ are available.
@@
6 Conclusion
We introduced HDGL, a novel and efficient algorithm for transductive learning on graphs using hyperdimensional representations. HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node. The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks. HDGL obviates the need for expensive iterative training of the sort required by deep learning methods. HDGL is also well-suited for class incremental learning where the model has to learn to effectively discriminate between a growing number of classes.
Our experiments with widely used benchmark datasets show that on the node classification task, HDGL is competitive with the state-of-the-art GNN methods with respect to accuracy, at substantially reduced computational cost. Our experiments also show that the HD representation constructed by HDGL supports link prediction at accuracies comparable to that of DeepWalk and related methods. We conclude that HDGL offers a computationally efficient alternative to graph neural networks for node classification, especially in settings that call for class-incremental learning or in applications that demand high accuracy models at significantly lower computational cost and learning time than possible with the state-of-the-art graph neural network methods.

### v3

@@ 5.5. Implementation Details
Table 3. Class Incremental Node Labeling Results. No hyperparameter grid search is used for GCN and SGC algorithms. Wall-time (in seconds) includes both training and inference of the model. HDGL does not require retraining from $t=2$.
@@
5.6. Results
Node Classification Performance. The results of our comparison of HDGL with the other methods are shown in Table 2. We observe that HDGL without the need for computationally expensive iterative training or hyperparameter optimization achieves node prediction accuracy that is competitive with those of state-of-the-art GNN models (GCN, GAT, SGC) across most of the benchmark data sets, in each case reaching performance that is within 1-2% of the best performing model. HDGL also consistently outperforming basic deep learning baseline methods (LogReg, DeepWalk, LabelProp). The notable advantage of HDGL over RelHD lies in its ability to accommodate features beyond the binary $\{0,1\}$ range, as evident from the results presented in Table 2, where HDGL significantly outperforms. However, when dealing with datasets featuring $\{0,1\}$ attributes like CORA and Citeseer, the performance of HDGL and RelHD appears nearly identical.
It is important to note that although we used grid search over 12 configurations for hyperparameter optimization, typically deep learning models explore a much larger parameter space. This expanded search naturally comes with increased runtime, especially when GPUs are utilized, further widening the energy efficiency gap between Graph Neural Networks (GNN) and traditional deep learning models like HDGL.
@@ 5.6. Results
Table 4. Link prediction results averaged over 10 Randomized Instantiations of HDGL and datasplits. Results for other models are from Kipf and Welling (2016) paper.
Methods Cora Citeseer PubMed
AUC AP AUC AP AUC AP
DeepWalk 0.831 0.850 0.805 0.836 0.844 0.841
SC 0.846 0.885 0.791 0.826 0.849 0.888
GAE 0.910 0.920 0.895 0.899 0.964 0.965
VGAE 0.914 0.926 0.908 0.920 0.944 0.947
HDGL 0.849 0.880 0.768 0.842 0.853 0.884
Figure 3. Link Prediction Performance of HDGL under Various Dimensionality Configurations.
[…]
HDGL in the Class-Incremental Learning Setting Consider a scenario with a dynamic graph where nodes and edges remain static, but the new node labels appear over time i.e initially, only a subset of nodes is labeled, belonging to two classes. Over time, new nodes receive labels; thus, introducing new classes. This is a node labeling instance of the class-incremental learning problem (Van de Ven et al., 2022; Belouadah et al., 2021; Tian et al., 2024). In this setting, deep learning methods typically require retraining on the data for previously learned classes along with the data for the new classes to avoid catastrophic forgetting of the previously learned classes (Tian et al., 2024). This retraining incurs significant additional runtime and computation.
@@
To simulate incremental node classification on a fully labeled graph, we use BlogCatalog and Coauthor Physics Dataset. We first split the nodes into train/val/test as in Shchur et al. (2018). We mask the train/val/test such that at time $t=1$ only the labels $\{1,2\}$ are available; and at $t=2$ only the labels $\{1,2,3\}$ are available; and so on such that at the final time step, all labels $\{1,\cdots,L\}$ are available.
@@
6. Conclusion
We introduced HDGL, a novel transductive learning algorithm for graphs using hyperdimensional representations. HDGL offers a computationally efficient alternative to graph neural networks by leveraging the injectivity property of node representations from Graph Neural Networks (GNNs). It uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node. The resulting latent node representations support both node classification and link prediction tasks. Furthermore, HDGL eliminates the need for iterative training, making it ideal for class-incremental learning and applications requiring high accuracy models at lower computational cost and learning time compared to traditional graph neural network methods. The advantages and efficacy of HDGL are validated by comprehensive experiments comparing it with state-of-the-art GNN methods.
Acknowledgements.
This work was funded in part by grants from the National Science Foundation (2226025), the National Center for Advancing Translational Sciences, and the National Institutes of Health (UL1 TR002014)